OpenStudy (anonymous):

Factor completely and show all work step by step: xy-2x+4y-8

OpenStudy (anonymous):

xy-2x+4y-8= x(y-2)+4(y-2)= (y-2)(x+4)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

no prob